A merchant has coffee worth $3 a pound that she wishes to mix with 6 pounds of coffee worth $8 a pound to get a mixture that can be sold for $4 a pound. How many pounds of the $3 coffee should be used?
Answer by Nate(3500)

You want an average of the two:
(Amount + Amount)/(Total Amount of Nuts) = Cost Average
Prices of each is included:
(Amount(Price) + Amount(Price))/(Total Amount of Nuts) = Cost Average
(6($8.00) + x($3.00))/(x + 6) = $4.00
48 + 3x = 4(x + 6) = 4x + 24
24 pounds = x